A bullet 2.00 cm long is fired at 420π‘š/𝑠 and passes straight through a 10.0 cm thick board, exiting at 280π‘š/𝑠. What is the average acceleration of the bullet through the board?


a. -490,000 m/ s2

b. 196,000 m/ s2

c. -1,400 m/ s2

d. -700 m/s2

1
Expert's answer
2022-02-10T09:18:04-0500

The average acceleration of the bullet through the board

"a=\frac{v_f^2-v_i^2}{2d}"

"a=\frac{280^2-420^2}{2*0.10}=-490\:000\:\rm m/s^2"

Answer: A